OpenStudy (monica0597):

I just need to know how to solve this. Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

y + 1 = 4(x + 3)

OpenStudy (anonymous):

y - y1 = m(x - x1)

OpenStudy (monica0597):

THANK yOU!

OpenStudy (anonymous):

you welcome! <3